Rooney Company began operations on January 1, 2018, by issuing common stock for $35,000 cash. During 2018, Rooney received $58,700 cash from revenue and incurred costs that required $38,700 of cash payments.

Prepare a GAAP-based income statement and balance sheet for Rooney Company for 2018, under each of the following independent scenarios:

  1. Rooney is a promoter of rock concerts. The $38,700 was paid to provide a rock concert that produced the revenue.

  1. Rooney is in the car rental business. The $38,700 was paid to purchase automobiles. The automobiles were purchased on January 1, 2018, and have three-year useful lives, with no expected salvage value. Rooney uses straight-line depreciation. The revenue was generated by leasing the automobiles.

  1. Rooney is a manufacturing company. The $38,700 was paid to purchase the following items:

  1. (1) Paid $3,300 cash to purchase materials that were used to make products during the year.

  2. (2) Paid $2,380 cash for wages of factory workers who made products during the year.

  3. (3) Paid $17,820 cash for salaries of sales and administrative employees.

  4. (4) Paid $15,200 cash to purchase manufacturing equipment. The equipment was used solely to make products. It had a three-year life and a $2,000 salvage value. The company uses straightline depreciation.

  5. (5) During 2014, Lang started and completed 2,100 units of product. The revenue was earned when Lang sold 1,700 units of product to its customers.
Answer #1

Part A

Rooney Company

Income Statement 2018

Sales revenue

58700

Operating expenses

(38700)

Net income (Loss)

$20000

Rooney Company

Balance Sheet as of 12/31/2018

Assets

Cash

55000

Total Assets

$55000

Equity

Common stock

35000

Retained earnings

20000

Total equity

$55000

35000+20000 = $55000

Retained earnings = net income

Part B

Rooney Company

Income Statement 2018

Sales revenue

58700

Depreciation expenses

(12900)

Net income (Loss)

$45800

38700/3 = 12900

Rooney Company

Balance Sheet as of 12/31/2018

Assets

Cash

55000

Rental equipment

38700

Accumulated depreciation

(12900)

Total Assets

$80800

Equity

Common stock

35000

Retained earnings

45800

Total equity

$80800

35000+20000 = $55000

Part C

Rooney Company

Income Statement 2018

Sales revenue

58700

Cost of goods sold

(8160)

Gross margin

50540

Administrative expense

(17820)

Net income (Loss)

$32720

Depreciation for manufacturing equipement = (15200-2000)/3 = $4400

Product cost = 3300+2380+4400 = 10080

Per unit product cost = 10080/2100 = $4.80

Cost of goods sold = 1700*4.80

Rooney Company

Balance Sheet as of 12/31/2018

Assets

Cash

55000

Finished goods inventory

1920

Manufacturing equipment

15200

Accumulated depreciation

(4400)

Total Assets

$67720

Equity

Common stock

35000

Retained earnings

32720

Total equity

$67720

35000+20000 = $55000

Finished goods inventory = (2100-1700)*4.80